# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

### v2

@@ Abstract
We introduce a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal of MDMS is to maximize an objective function combining a monotone submodular utility term and a min-distance diversity term between any pair of selected points, subject to a cardinality constraint. We propose the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove that it is NP-hard to approximate to within a factor of $0.5584$. Finally, we demonstrate that GIST outperforms existing benchmarks for on a real-world image classification task that studies single-shot subset selection for ImageNet.
@@ 1 Introduction
Subset selection is a challenging combinatorial optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, drug discovery, data summarization, and designing pretraining sets for large language models (Team et al., 2024). Data sampling in particular is a salient problem due to unprecedented and continuous streams of data collection. For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
@@
In many subset selection tasks, we rely on the utility (or weight) of an object to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects. If we select a small subset, then we must ensure that the selected set is a good representation of the original set. These utility, diversity, and coverage criteria are expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees. The algorithmic tools used in constrained subset selection shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
@@
This work introduces a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS). Given $n$ points $V$ in a metric space, a monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, and a cardinality constraint $k$, the goal is to solve
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}g(S)+\texttt{div}(S)$ (1)
subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k,$
where the diversity term $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the min distance between any two distinct points in the selected set. Meinl et al. (2011) call $\texttt{div}(S)$ the $p$-dispersion measure, and Bhaskara et al. (2016) call it the min-min diversity objective, where is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
Other diversity terms such as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016; Borodin et al., 2017). While previous works have explored maximizing $\texttt{div}(S)$ in isolation, MDMS introduces a utility term and has not been studied before. Bhaskara et al. (2016) remark that $\texttt{div}(S)$ is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point. Our work shows combining a monotone submodular utility term with $\texttt{div}(S)$ mitigates some of the challenges associated with this diversity measure, making it easier to leverage its properties.
@@ 1.1 Our Contributions
We summarize our main contributions below:
  • In Section 2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up, demonstrating how the competing terms in the objective function interact.
  • In Section 3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution. In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
[…]
  • In Section 5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm). Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
@@ 1.2 Related Work
Diversity objectives.
A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its nearest center in $S$. Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018). The main difference between MDMS and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting. The second key difference is that we relax the requirement to cover every point in the dataset, which can be helpful for large datasets where we want to avoid instability due to certain corner or outlier points.
Bhaskara et al. (2016) study various distance-based diversity metrics subject to a cardinality constraint. In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation. They complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption. They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, but they enforce a strict cardinality constraint $|S|=k$; otherwise, it is optimal to only output one or two points. Meinl et al. (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization. This exact cardinality issue does not arise in MDMS because we need to balance the trade-off between competing monotone submodular utility and diversity terms.
Moumoulidou et al. (2020), Addanki et al. (2022), and Mahabadi and Trajanovski (2023) study diversity maximization under fairness constraints. Given a partitioning of the dataset, they ensure that a specified number of points $k_{i}$ are selected from each group $i\in[m]$. This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$). Moumoulidou et al. (2020); Addanki et al. (2022) maximize $\texttt{div}(S)$ in this setting. Mahabadi and Trajanovski (2023) consider sum-of-all pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives. However, none of these works add a utility term to the objective.
Borodin et al. (2017) optimize the combination of a submodular utility and sum-of-all pairwise distances in the selected set, making it the most similar work to MDMS. Unlike our $\texttt{div}(S)$ term, max-sum diversity is not very sensitive to duplicate points. In fact, adding $m$ copies of the same point to a set raises the increment of the diversity term by a factor of $m$ in their model. By using $\texttt{div}(S)$ instead, we can strictly avoid duplication across all selected points.
@@
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Ash et al. (2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity. Wei et al. (2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. (2021) cluster examples that are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. (2019); Zhdanov (2019)) in that we directly incorporate the utility and diversity terms in our objective function.
@@ 2 Preliminaries
Here we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a starting point.
Submodular set function.
For a set function $g:2^{V}\rightarrow\mathbb{R}$ and any $S,T\subseteq V$, let $g(T\mid S)=g(S\cup T)-g(T)$ denote the marginal gain of $g$ at $T$ with respect to $S$. A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singleton sets. A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
[…]
Min-distance diversity.
For a set of $n$ points $V$ in a metric space, define the min-distance diversity function
$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
For $\texttt{div}(S)$ to be well-defined, we set the diversity to be the diameter of $V$ if $|S|\leq 1$, i.e., $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$. This implies that it is always beneficial to select at least two points to maximize the overall objective.
We extend the distance function to take subsets $S\subseteq V$ as input, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$. For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
Optimization problem.
Define the objective function
$f(S)=g(S)+\texttt{div}(S).$ (2)
The goal of the min-distance diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
@@
The objective function $f(S)$ is not submodular. We present a counterexample in Appendix A.
@@
Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges
$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
@@
The objective function $f(S)$ has two terms, each of which is easy to (approximately) optimize in isolation. For the $g(S)$ term, run the greedy algorithm to get $S_{1}$ satisfying
@@
For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$. The algorithm then flips a biased coin and outputs $S_{1}$ with probability $p\approx 0.613$, and $S_{2}$ otherwise. This achieves a $0.387$-approximation guarantee:
$\mathbb{E}[\textnormal{ALG}_{\textnormal{simple}}]$ $\geq p\cdot g(S_{1})+(1-p)\cdot\texttt{div}(S_{2})$
$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\texttt{div}(S^{*})$
@@
where we take $p=e/(2e-1)$ by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of the main goals of this paper is to improve over this baseline and prove complementary hardness results.
@@ 3 Algorithm
In this section, we present the GIST algorithm and prove that it achieves an approximation ratio of $(\nicefrac{{1}}{{2}}-\varepsilon)$ for MDMS. GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine to find a high-value independent set of the intersection graph $G_{d}(V)$. This is in contrast to the simple $0.387$-approximation algorithm that just tries two extreme thresholds $D=\{0,d_{\max}\}$.
The GreedyIndependentSet heuristic builds a set $S$ (starting from the empty set) by repeatedly adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular function $g$ while satisfying $\textnormal{dist}(v,S)\geq d$. This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
[…]
Algorithm 1 Diverse data summarization for submodular utilities using greedy-weighted maximal independent sets.
1:function GIST(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$triangleright runs the classic greedy algorithm
@@
7:     Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$triangleright distance thresholds
@@
1:function GreedyIndependentSet(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
@@
6:              return $S$triangleright $S$ is a maximal independent set of $G_{d}(V)$
@@
For any $\varepsilon>0$, the GIST algorithm outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers and adapted for monotone submodular rewards.
@@
[…]
@@
We first construct an injective map $h:S_{d}^{*}\rightarrow T$. We say that a point $v\in V$ is covered by $B_{i}$ if $v\in B_{i}$. For each covered $s\in S_{d}^{*}$, we map it to $t_{i}$, where $i$ is the minimum index for which $B_{i}$ covers $s$. If there is an $s\in S_{d}^{*}$ not covered by any set $B_{i}$, then GreedyIndependentSet must have selected $k$ points, i.e., $|T|=k$. We map the uncovered points in $S_{d}^{*}$ to arbitrary points in $T$ while preserving the injective property.
@@
We use $h$ to account for $g(s\mid T)$ in terms of the values we gained in set $T$. For any $s\in S_{d}^{*}$, let $T_{s}$ be the set of points added to set $T$ in algorithm GreedyIndependentSet right before the addition of $h(s)$. Since GreedyIndependentSet incrementally adds points and never deletes points from $T$, we know $T_{s}\subseteq T$. By submodularity, $g(s\mid T)\leq g(s\mid T_{s})$. By the greedy nature of the algorithm, we also know that
@@
Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of GIST.
Proof of Theorem 3.1.
Let $d^{*}$ be the minimum pairwise distance between points in $S^{*}$. The GIST algorithm iterates over a set of thresholds $D$. The definition of $D$ implies that at least one threshold $d$ is in the range $[d^{*}/(2(1+\varepsilon)),d^{*}/2)$ unless $d^{*}\leq\varepsilon d_{\max}$. We deal with this special case later and focus on the case when such a $d$ exists.
Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Applying Lemma 3.2 since $d<d^{*}/2$, we know that
@@
It remains to prove the claim for the special case of $d^{*}\leq\varepsilon d_{\max}$. Algorithm GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$. The first set is the classic greedy solution for the submodular function $g$, ignoring the diversity term. This gives the lower bound (Nemhauser et al., 1978):
$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\varepsilon d_{\max}\right).$
The other initial set realizes the max pairwise distance $d_{\max}$ and ignores the submodular term, yielding the lower bound:
$f(T)\geq d_{\max}.$
Combining these two lower bounds, we get a final bound of
$f(T)\geq\left(1-\frac{1}{e}\right)\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$
which completes the proof. ∎
Remark 3.3.
We can modify GIST to guarantee an exact $\nicefrac{{1}}{{2}}$-approximation ratio by trying all pairwise thresholds $D^{\prime}=\{\textnormal{dist}(u,v)/2-\varepsilon^{\prime}:u,v\in V\}$, for a small enough $\varepsilon^{\prime}>0$, at the cost of $O(n^{2})$ calls to GreedyIndependentSet instead of $O(\log_{1+\varepsilon}(1/\varepsilon))$.

### v3

# GIST: Greedy Independent Set Thresholding for Max-Min Diversification with Submodular Utility

@@ Abstract
This work studies a novel subset selection problem called max-min diversification with monotone submodular utility (MDMS), which has a wide range of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal of MDMS is to maximize $f(S)=g(S)+\lambda\cdot\texttt{div}(S)$ subject to a cardinality constraint $|S|\leq k$, where $g(S)$ is a monotone submodular function and $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ is the max-min diversity objective. We propose the GIST algorithm, which gives a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove that it is NP-hard to approximate within a factor of $0.5584$. Finally, we show in our empirical study that GIST outperforms state-of-the-art benchmarks for a single-shot data sampling task on ImageNet.
@@ 1 Introduction
Subset selection is a ubiquitous and challenging problem at the intersection of machine learning and combinatorial optimization, finding applications in areas such as feature selection, recommendation systems, and data summarization, including the critical task of designing pretraining sets for large language models [59, 3]. Data sampling in particular is an increasingly important problem given the unprecedented and continuous streams of data collection. For example, one self-driving vehicle generates ~80 terabytes of data daily from LiDAR and imaging devices [34], and academic datasets have scaled dramatically from 1.2M images in ImageNet [54] to 5B in Laion [56].
@@
Subset selection often involves balancing competing objectives: we rely on the utility (or weight) of individual items to prioritize them, while simultaneously trying to avoid selecting duplicate or near-duplicate items to ensure diversity. When selecting a small subset, this process should guarantee that the chosen set is a good representation of the original dataset—often called coverage. The intricate trade-offs between utility, diversity, and coverage are often expressed through an objective function. It is often a significant challenge to design efficient algorithms with strong approximation guarantees for constrained subset selection problems, even when leveraging tools and techniques from combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
@@
This paper studies a novel subset selection problem called max-min diversification with monotone submodular utility (MDMS). Given $n$ points $V$ in a metric space, a nonnegative monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, diversity strength $\lambda\geq 0$, and cardinality constraint $k$, our goal is to solve
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~g(S)+\lambda\cdot\texttt{div}(S)$ (1)
$\hskip 13.6572pt\text{subject to}~~|S|\leq k,$
where $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ is the max-min diversification term [2, 40, 38]. Maximizing the submodular term $g(S)$ aims to select points that are the most valuable or informative. A purely utility-driven algorithm, however, can lead to redundancies where very similar points are chosen. To counteract this, we add the $\lambda\cdot\texttt{div}(S)$ term to encourage selecting points that are well spread out in the metric space, which effectively penalizes subsets with closely clustered elements. Note that $\lambda$ is a knob that controls the trade-off between these two terms, acting as a regularization strength for subset selection problems.
The most relevant prior work is by Borodin et al. [13]. They combine a monotone submodular utility and the max-sum diversity term $\sum_{u,v\in S}\textnormal{dist}(u,v)$, in contrast to our max-min $\texttt{div}(S)$ in (1). They give a greedy $\nicefrac{{1}}{{2}}$-approximation algorithm and extend their results to general matroid constraints via local search. An earlier contribution by Gollapudi and Sharma [26] combined a linear utility (instead of a monotone submodular function) and the max-sum diversity term, subject to the strict constraint $|S|=k$, to design a ranking system for search engines. These approaches focus on maximizing the sum of pairwise distances, but our max-min formulation ensures that the selected points have high utility and are maximally separated, leading to less redundant and more representative subsets.
@@ 1.1 Our contributions
We summarize the main contributions of this work below:
  • In Section˜2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up to show how the competing terms in the objective function interact.
  • In Section˜3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution. In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
  • In Section˜4, we prove that it is NP-hard to approximate MDMS to within a factor of $0.5584$ via a careful reduction from the maximum coverage problem. For linear utilities, we match the guarantees of GIST and prove a tight $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, assuming $\text{P}\neq\text{NP}$. In the more restricted case of a linear utility function and the Euclidean metric, we prove APX-completeness.
  • In Section˜5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm). Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
@@ 1.2 Related work
Submodular maximization.
Submodular maximization subject to a cardinality constraint $k$ is an NP-hard problem, i.e., $\max_{S\subseteq V:|S|\leq k}g(S)$. For monotone submodular functions, Nemhauser et al. [47] proved that the greedy algorithm achieves a $(1-\nicefrac{{1}}{{e}})$-approximation, which is asymptotically optimal unless $\text{P}=\text{NP}$ [23]. The non-monotone case is substantially less understood, but Buchbinder and Feldman [14] recently gave a 0.401-approximation algorithm and Qi [48] improved the hardness of approximation to 0.478. Over the last decade, there has also been significant research on distributed algorithms for submodular maximization in the MapReduce [44, 43, 9, 10, 39, 33] and low-adaptivity models [7, 19, 22, 8, 5, 16, 17].
Diversity maximization.
The related max-min diversification problem $\max_{S\subseteq V:|S|=k}\texttt{div}(S)$ has a rich history in operations research due to its connection to facility location, and is also called the $p$-dispersion problem. It uses the strict constraint $|S|=k$; otherwise, if $|S|\leq k$ we can maximize $\texttt{div}(S)$ with two diametrical points. Tamir [58] gave a simple greedy $\nicefrac{{1}}{{2}}$-approximation algorithm for max-min diversification, and Ravi et al. [51] proved $(\nicefrac{{1}}{{2}}+\varepsilon)$-inapproximability, for any $\varepsilon>0$, unless $\text{P}=\text{NP}$. Indyk et al. [30] gave a distributed $\nicefrac{{1}}{{3}}$-approximation algorithm via composable coresets for diversity and coverage maximization. Borassi et al. [12] designed a $\nicefrac{{1}}{{5}}$-approximation algorithm in the sliding window model. For asymmetric distances, Kumpulainen et al. [38] recently designed a $1/(6k)$-approximation algorithm. There are also many related works on mixed-integer programming formulations, heuristics, and applications to drug discovery [20, 52, 55, 60, 37, 40, 61, 42].
Bhaskara et al. [11] studied the sum-min diversity problem $\max_{S\subseteq V:|S|\leq k}\sum_{u\in S}\textnormal{dist}(u,S\setminus\{u\})$, providing the first constant-factor approximation algorithm with a $\nicefrac{{1}}{{8}}$ guarantee. Their algorithm is based on a novel linear-programming relaxation and generalizes to matroid constraints. They also proved the first inapproximability result of $\nicefrac{{1}}{{2}}$ under the planted clique assumption.
Diversity maximization has also recently been studied with fairness constraints. Given a partition of the points into $m$ groups, they ensure $k_{i}\in[\ell_{i},u_{i}]$ points are selected from each group $i\in[m]$, subject to $|S|=k=\sum_{i=1}^{m}k_{i}$. This line of work can be categorized as studying fair max-min [46, 2, 61], sum-min [11, 41], and max-sum [1, 41] diversity objectives.
@@
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Greedy coreset methods based on $k$-center have been successfully used for data selection with and without utility terms [50, 57]. Ash et al. [6] use $k$-means++ seeding over the gradient space of a model to balance uncertainty and diversity. Wei et al. [62] introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. [18] cluster examples represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. [35], Zhdanov [63]) in that we directly incorporate the utility and diversity terms into the objective function.
@@ 2 Preliminaries
Submodular function.
For any $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$ and $S,T\subseteq V$, let $g(S\mid T)=g(S\cup T)-g(T)$ be the marginal gain of $g$ at $S$ with respect to $T$. A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singletons. A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
Max-min diversity.
For a set of $n$ points $V$ in a metric space, define the max-min diversity function as
$\texttt{div}(S)$ $=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
We take $\texttt{div}(S)$ to be the diameter of $V$ if $|S|\leq 1$ so that it is monotone decreasing. We extend the distance function to take subsets $S\subseteq V$ as input in the standard way, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$, and define $\textnormal{dist}(u,\varnothing)=\infty$.
MDMS problem statement.
For any nonnegative monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$ and $\lambda\geq 0$, let
$f(S)=g(S)+\lambda\cdot\texttt{div}(S).$ (2)
The max-min diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
@@
The objective function $f(S)$ is not submodular (see Appendix˜A for a counterexample).
@@
Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges $E=\{(u,v)\in V^{2}:u\neq v,\textnormal{dist}(u,v)<d\}.$ For any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
@@
The objective function $f(S)$ is composed of two terms, each of which is easy to (approximately) optimize in isolation. For the submodular term $g(S)$, run the greedy algorithm to get $S_{1}$ satisfying
@@
For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$. Then, return the better of the two solutions. This gives a $0.387$-approximation guarantee because, for any $0\leq p\leq 1$, we have:
$\textnormal{ALG}_{\textnormal{simple}}$ $=\max\left\{f(S_{1}),f(S_{2})\right\}$
$\geq p\cdot g(S_{1})+(1-p)\cdot\lambda\cdot\texttt{div}(S_{2})$
$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\lambda\cdot\texttt{div}(S^{*})$
@@
where we set p=e/(2e1)p=e/(2e-1) at the end by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of our main goals is to improve over this baseline and prove complementary hardness of approximation results.
One of the most common algorithms for subset selection problems is the greedy algorithm that iteratively adds the item with maximum marginal value with respect to the objective function $f$. We show in Appendix B that the greedy algorithm does not provide a constant-factor approximation guarantee.
@@ 3 Algorithm
In this section, we present the GIST algorithm for the MDMS problem and prove that it achieves an approximation ratio of $\nicefrac{{1}}{{2}}-\varepsilon$. At a high level, GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine on the intersection graph $G_{d}(V)$ to find a high-valued maximal independent set. This is in contrast to the warm-up $0.387$-approximation algorithm, which only considers the two extreme thresholds $d\in\{0,d_{\max}\}$.
GreedyIndependentSet builds the set $S$ (starting from the empty set) by iteratively adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular utility $g$ while satisfying $\textnormal{dist}(v,S)\geq d$. This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
GIST computes multiple solutions and returns the one with maximum value $f(S)$. It first runs the classic greedy algorithm for monotone submodular functions to get an initial solution, which is equivalent to calling GreedyIndependentSet with distance threshold $d=0$. Then, it considers the set of distance thresholds $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$. The set $D$ contains a threshold close to the target $d^{*}/2$ where $d^{*}=\texttt{div}(S^{*})$. [^1] If this is not true, we are in the special case where $d^{*}\leq\varepsilon d_{\max}$, which we analyze separately. For each $d\in D$, GIST calls $\textnormal{{GreedyIndependentSet}}(V,g,d,k)$ to find a set $T$ of size at most $k$. If $T$ has a larger objective value than the best solution so far, it updates $S\leftarrow T$. After iterating over all thresholds in $D$, it returns the highest-value solution among all candidates.
Algorithm 1 Max-min diversification with submodular utility via greedy weighted independent sets.
1:function GIST(points $V$, monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:  Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$ $\triangleright$ classic greedy algorithm
@@
7:  Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$ $\triangleright$ distance thresholds
@@
1:function GreedyIndependentSet(points $V$, monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
@@
6:     return $S$ $\triangleright$ $S$ is a maximal independent set of $G_{d}(V)$
@@
For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$ using $O(nk\log_{1+\varepsilon}(1/\varepsilon))$ submodular value oracle queries.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-center and adapted for monotone submodular utility functions.
@@
Let $T$ be the output of $\textnormal{{GreedyIndependentSet}}(V,g,d^{\prime},k)$. If $d^{\prime}<d/2$, then $g(T)\geq g(S_{d}^{*})/2$.
@@
First we construct an injective map $h:S_{d}^{*}\rightarrow T$. We say that $B_{i}$ covers a point $v\in V$ if $v\in B_{i}$. For each covered $s\in S_{d}^{*}$, we map it to $t_{i}$, where $i$ is the minimum index for which $B_{i}$ covers $s$. If there is an $s\in S_{d}^{*}$ not covered by any set $B_{i}$, then GreedyIndependentSet must have selected $k$ points, i.e., $|T|=k$. We map the uncovered points in $S_{d}^{*}$ to arbitrary points in $T$ while preserving the injective property.
@@
We use $h$ to account for $g(s\mid T)$ in terms of the values we gained in set $T$. For any $s\in S_{d}^{*}$, let $T_{s}$ be the set of points added to set $T$ in algorithm GreedyIndependentSet right before the addition of $h(s)$. Since GreedyIndependentSet iteratively adds points and never deletes points from $T$, we know $T_{s}\subseteq T$. By submodularity, $g(s\mid T)\leq g(s\mid T_{s})$. By the greedy nature of the algorithm, we also know that
@@
Now that we have with this bicriteria approximation, we can analyze the approximation ratio of GIST.
Proof of Theorem˜3.1.
We first prove the oracle complexity of GIST. There are $1+\log_{1+\varepsilon}(2/\varepsilon)=O(\log_{1+\varepsilon}(1/\varepsilon))$ thresholds in $D$. For each threshold, we call GreedyIndependentSet once. In each call for $k$ iterations, we find the maximum marginal-value point by scanning (in the worst case) all points. This requires at most $nk$ oracle calls yielding the overall oracle complexity of the algorithm.
Let $d^{*}$ be the minimum pairwise distance between points in $S^{*}$. The GIST algorithm iterates over a set of thresholds $D$. The definition of $D$ implies that at least one threshold $d$ is in the interval $d^{*}/(2(1+\varepsilon)),d^{*}/2)$, unless $d^{*}\leq\varepsilon d_{\max}$. We deal with this special case later and focus on the case when such a $d$ exists.
Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Since $d<d^{*}/2$, Lemma˜3.2 implies that
@@
It remains to prove the claim for the case when $d^{*}\leq\varepsilon d_{\max}$. GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$. The first set is the classic greedy solution [47] for the monotone submodular function $g(S)$, and ignores the diversity term. It follows that
$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\lambda\cdot\varepsilon d_{\max}\right).$ (3)
The second set contains two points with pairwise distance $d_{\max}$, and ignores the submodular term. This yields the lower bound
$f(T)\geq\lambda\cdot d_{\max}\implies-\lambda\cdot d_{\max}\geq-f(T).$ (4)
Combining (3) and (4), we get a final bound of $f(T)\geq(1-\frac{1}{e})\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$ which completes the proof. ∎